OpenStudy (ray_sihota):

how would I graph 3y+1=7

OpenStudy (ray_sihota):

having issue graphing 3x-4y-5=-17

OpenStudy (ray_sihota):

as well

OpenStudy (jdoe0001):

solve for "y" on both cases, then pick a couple of random "x" values and get "y", and plot it

OpenStudy (ray_sihota):

3y+1=7 =3y=6 y=2

OpenStudy (ray_sihota):

so would that make x =1

OpenStudy (jdoe0001):

yeap so y =2 that means \(\large \begin{array}{rrllll} x&y \\\hline\\ \pm 5&2\\ \pm 50&2\\ \pm 5,000&2\\ \pm 5,000,000&2\\ \pm \infty&2 \end{array}\) thus 

OpenStudy (ray_sihota):

so the 3x-4y-5=-17 would that have be made into the y=mx+b form 4y=-3x+4y=-17 so then for the 3y+1=7 eqation do not have to find the other two points.

OpenStudy (jdoe0001):

right.. the \(\bf 3y+1=7 \) is just a horizontal line... you could look for two points, look above... there are more than two.. then again.. they'll all end up over the y =2 line anyhow as far as the 2nd one \(\bf 3x-4y-5=-17\implies -4y-5=-17-3x\implies -4y=-17-3x+5 \\ \quad \\ -4y=-12-3x\implies y=\cfrac{-12-3x}{-4}\implies y=\cfrac{\cancel{ -12 }}{\cancel{ -4 }}-\cfrac{-3x}{-4} \\ \quad \\ y=3+\cfrac{3x}{4}\)

OpenStudy (jdoe0001):

ok so it got a bit truncated... so \(\bf 3x-4y-5=-17\implies -4y-5=-17-3x \\ \quad \\ -4y=-17-3x+5 \implies-4y=-12-3x\implies y=\cfrac{-12-3x}{-4} \\ \quad \\ y=\cfrac{\cancel{ -12 }}{\cancel{ -4 }}-\cfrac{-3x}{-4}\implies y=3+\cfrac{3x}{4}\)

OpenStudy (ray_sihota):

and then just substitute then into one of the equations to find x

OpenStudy (jdoe0001):

well... then just pick 2 random "x" and get "y" that would give you 2 coordinate points and since is a LINEar equation, its graph is a line, and to graph a line all you need is two points... .draw a line through those points

OpenStudy (ray_sihota):

got it thanks a lot for your help

OpenStudy (jdoe0001):

yw